# Compact LED-Based Displacement Sensing for Robot Fingers

### v2

@@ Abstract
In this paper, we introduce a sensor designed for integration in robot fingers, where it can provide information on the displacements induced by external contact. Our sensor uses LEDs to sense the displacement between two plates connected by a transparent elastomer; when a force is applied to the finger, the elastomer displaces and the LED signals change. We show that using LEDs as both light emitters an receivers in this context provides high sensitivity, allowing such an emitter and receiver pairs to detect very small displacements. We characterize the standalone performance of the sensor by testing the ability of a supervised learning model to predict complete force and torque data from its raw signals, and obtain a mean error between 0.05 and 0.07 N across the three directions of force applied to the finger. Our method allows for finger-size packaging with no amplification electronics, low cost manufacturing, easy integration into a complete hand, and high overload shear forces and bending torques, suggesting future applicability to complete manipulation tasks.
@@
Contact Sensing, LED emitter and receiver, Robot Hand, Robotic Manipulation
@@ I Introduction
The most direct method for sensing such forces and torques is to add an off-the-shelf six-axis force and torque (F/T) sensor into the base of the finger. However, such F/T sensors are primarily developed with industrial applications in mind [6, 7, 8], leading to limitations when being used for dexterous manipulation. Manipulation often benefits from compliance [9], whereas most F/T sensors are optimized to be as stiff as possible, thus requiring precision engineering and leading to a high cost. Industrial applications also tend to favor a higher resolution than is needed for manipulation, but compromise by limiting the overload protection [6]. This leads to sensors that are easily damaged in the uncertain environments that robotics need to interact with. Finally, off-the-shelf F/T sensors can be difficult to package and integrate into the hand. Others have attempted to solve this problem by designing sensors to be placed within the fingertips [10, 11, 12]. This limits the geometry of the finger and can also limit the inclusion of other modalities within the finger.
@@
In this paper, we present the design of a displacement sensor intended to provide information about the net forces acting on a robot finger in a manner that addresses the needs of robotic manipulation. First, such a sensor must be easy to manufacture using low-cost components and processes, so that it is affordable and practical to include multiple in a multi-fingered hand. Second, it must have the size and profile suited for the base of a fingertip, which will increase ease of integration into a hand. This also frees up the finger design for alternate sensors, or allows more flexibility in design of a finger. Third, some amount of compliance could be beneficial to manipulation, and thus there is no need to avoid compliance at all cost. Finally, the ideal sensor would have a Signal to Noise Ratio (SNR) that allows it to be useful for manipulation without sacrificing overload protection.
@@
Figure 1: Our LED-based sensor, standalone (left) and mounted to a fixed base and robot finger (right). We use LEDs mounted on two plates connected by a transparent elastomer to sense displacement between the plates and relate it to applied forces and torques. This method allows for a compact, fully integrated package at a low manufacturing cost.
A key feature of our work is that using LEDs as both emitters and receivers confers very high signal-to-noise ratio, significantly higher than using photodiodes. Combined with the small LED form factor, this allows us to sense very small displacements in a compact package. Using an off-the-shelf F/T sensor as ground truth, we show that our sensor has an observed resolution of .06 N and 2.6 N-mm in a compact, low-cost package that is compliant and easy to integrate in robot fingers as it requires no external amplification. To the best of our knowledge, this is the first time that LEDs have been integrated into a 6-DOF flexure, allowing for full contact information in a compact, complaint robot fingers. Our sensor is designed to be integrated at the base of the finger, which allows for sensing coverage of the entire finger and the ability to integrate other sensors in the finger. We hope that this direction can facilitate truly ubiquitous sensing integrated in robot fingers, with future applications in manipulation.
Figure 2: A diagram of our LED-to-LED sensor concept. We align two PCBs with LEDs directly above each other. We attach these circuit boards to rigid plates, we then suspend both of these plates in a translucent, elastic material. Assuming the bottom plate is fixed, a force applied to the top plate will displace the plate and change how to emitter LED shines on the receiving LEDs. We can then map this change to the contact wrench on the sensor that induced the displacement.
II Related Work
Many tactile sensors use LEDs due to their compact packaging and wide range of emission cones [13, 14, 15, 16, 17]. The LEDs in these sensors are embedded in a translucent elastomer, so that their emission cone shines on a receiving device. This receiver is often a photodiode [14, 16] or camera [15, 17, 18], but can also be an LED [19]. Regardless of the receiving device, when a force is applied to the surface of these sensors, the elastomer deforms and the emitter and receiver will displace. Using either a data-driven approach or deriving an analytical relationship, the change in electrical signal from this deformation can then be mapped to the contact applied. These tactile sensors are not able to cover the full surface of a robot finger, and therefore do not provide full sensing coverage all possible contacts. To detect the full contact, others have tried several modalities including strain-based transducers [20], capacitance [21], optical [22], and magnetic [12]. Although most of these sensors are too large to integrate in a robot finger, there are a few novel F/T sensors that are built particularly for manipulation.
@@ II Related Work
Figure 3: Graphs showing the sensitivity of LEDs and Photodiodes when used as light receivers. Left: This graph shows the sharper cone response of the LED when compared to the photodiode. Right: This graph is a zoom in on the ADC signal received across a .1mm window. It shows the clearer SNR when using the LEDs as receivers versus the dedicated photodiodes. Because the signal is spread over a smaller area, there are sharper changes at small displacements, including in a .1 mm window.
@@ III-A Concept
Our conceptual diagram of our sensor is shown in Fig. 2. It consists of two rigid plates connected by an elastomeric layer that allows 6-DOF displacements. The electronics fixed to each rigid plate consist of a circuit boards with emitter LEDs and clusters of receiving LEDs. At the neutral position, the emitter LED is directly above the center of the receiving cluster. When an external force is applied, the plates will displace relative to each other. This displacement will result in a change in signal for the receiver LEDs due to the relative motion of the corresponding emitter LED. We can then map this change in signal back to the contact wrench applied.
@@ III-B Analysis of Single Emitter-Receiver Pair
Displacement transduction via light measurements offers three primary advantages. First, it provides a fast response time, as diode readings are gated by the ADC rather than the diodes. Second, its components require minimal space, which aids with reducing our footprint. Third, we can select LEDs in the IR spectrum, making the sensor resistant to the majority of external environment light. Adding an opaque wrapping to the outside of our sensor, we can further prevent any external light interference from external infrared light.
Using LEDs as both a receiver and emitter gives us several advantages over the other modalities mentioned in Table I. First, light can be sensed at very high frequencies with minimal electrical noise. There are several devices we could use as our receiver, but we opted to use LEDs as the receivers. In our case, the decision of using receiving LEDs over photodiodes was made based on 3 factors. 1) LEDs have a smaller form factor than most photodiodes in the price range. 2) LEDs are optimized for operation in their wavelength range, making it easier to design features to minimize external interference. 3) LEDs have a smaller viewing angle, which is useful for increasing sensitivity given our size constraints.
We first verify the validity of LEDs as both emitters and receivers, and compare to the alternative using photodiodes on the receiving end. To that end, we measured the response to relative horizontal displacements of a single emitter-received pair separated by an air gap, where the receiver consisted of either and LED or a photodiode. The results, plotting receiver signal vs. relative horizontal displacement of the emitter, are shown in Fig. 3. We observe that the LED receiver exhibits much higher sensitivity compared to its photodiode counterpart. In fact, the LED receiver can identify displacements as small as 0.01 mm, whereas the corresponding signal changes for a photodiode receiver are comfortably inside the noise level. Based on this result, we proceed with LED-LED pairs for the overall sensor design.
Figure 4: Our sensor is composed of two custom PCBs that have both LED emitters and clusters of LED receivers. When the top and bottom PCB boards are aligned facing each other, our sensor has 3 total emitter-receiver pars. The PCBs are fastened to rigid, 3-D printed plates. The plates are then fixed above each other in a mold that we then fill with PDMS and cure to get the sensor on the right. We use a single JST connector to communicate between the top and bottom board. We use an additional JST connector on the bottom board to communicate with external electronics. The rigid plates are fitted with threaded holes to mount external hardware. Our sensor is 27 mm in diameter, 20mm in height, and weighs 16.4g.
@@ III-C Complete Sensor Design
With the core concept of using identical LEDs as emitters and receivers verified, we implemented a preliminary design for the sensor. As we wanteded to mount this at the base of a finger, we constrained our design to a 1 inch (25.4 mm) diameter. The goal was to maximize the amount of data signals within the space. Based on our LED analysis, we used a compact, infrared LED (SFH 4056) with an emission cone of $20^{\circ}$. The primary size constraint became the op-amps, which meant we could support a maximum of 12 signals per board. As shown in the left of Fig. 4, we divided the signals into clusters of 4 receivers, so that each board had 3 clusters, for a total of 24 signals broken across 12 clusters. We then suspended the PCB boards above each other and attached them with an elastic flexure.
We fastened the bottom and top PCB boards, as shown in Fig. 4, each to a rigid, 3-D printed plate. The plates each have three M2 inserts that we can mount external hardware too. We used these inserts to attach these rigid plates to a mold that we then filled with an uncured mixture of PDMS. We cured the filled mold at $70^{\circ}C$ for 8 hours, which gave us the sensor shown in Fig. 4. The overall dimensions of the sensor is 20.82mm in height and a diameter of 27mm, which is just above our target diameter. We then covered the sensor with a thin, opaque rubber shrink-wrap tube. This cover inhibits interference from external light sources, but also helps adds to the sensor robustness. We cured the PDMS at a 10:1 ratio of base to curing agent, as we observed this best achieved the desired sensor stiffness and surface finish. Moreover, this ratio produced the most consistent behavior across multiple sensors.
@@
Figure 5: Mechanical hysteresis testing over three consecutive loading and unloading cycles. We fixed our sensor, with its rubber cover, on a reference F/T sensor and then used a linear probe to displace the top plate by 1 mm. Plotting the force and torque vs displacement for three consecutive cycles of pokes, we observe some hysteresis that is relatively consistent between cycles.
Figure 6: Mechanical overload forces and torques for our sensor. We built five sensors and then used an ADMET instron machine to apply a horizontal force until the sensor failed. By fixing it to the machine in different ways, we tested our sensor in five major failure modes - shear, bending, torsion, tension, compression. All tests were done with the opaque, shrink wrapped rubber cover on. Our sensor does not break under very high forces and torques, which is ideal for contact rich manipulation in uncertain robot environments.
@@ IV Mechanical Characterization
A core design objective of this sensor is robustness against large displacements induced by high contact forces. We tested our sensor to failure in five failure modes - shear, bending, torsion, tension, and compression. We note that, since our sensor is intended to be mounted at the base of a finger, as shown in Fig. 1, contact forces will mostly cause shear and bending of our sensor. Additionally, all tests on our sensor were done with a shrink wrapped neoprene cover, as shown on the finger in Fig. 1. This cover improved the robustness of our sensor, as it added an extra layer of protection to the PDMS flexure, especially for shear and tension forces.
The magnitude and direction of the forces and moments for each failure mode relative to our sensor are shown in Fig. 6. In particular, we note that our sensor withstands shear forces of up to 110 N and bending torques of up to 1.62 N*m. The overload forces for torsion and compression are also high. Our sensor is weakest in tension, which is not a major concern as a robot finger would likely not experience large tension forces while doing manipulation in uncertain environments.
PDMS is a hysteric material, so we expect our 6-DOF flexure to display some time-dependent non-linearity [13, 25]. In Fig. 5, we plot the results of cyclic loading and unloading tests for our sensor. We used a linear probe to horizontally displace the top plate of our sensor while it was fixed to a reference, off-the-shelf F/T sensor. Displacing the top plate by 1 mm three consecutive times, we observe hysteresis that is relatively consistent over the cycles. While hysteresis is a common effect for sensors that use PDMS as an elastic flexure [25], it is also problematic for mapping the electrical signals to the force applied, and often displays behavior that is unique to each sensor. In future work, we aim to investigate different clear elastomers for reducing this effect, potentially in combination with more complex learning architectures that can aim to learn this behavior and reduce in post processing.
@@ V Performance
In this section, we aim to quantify the performance of our design as a standalone sensor. Ultimately, we believe that the true value of a sensor designed for robot fingers lies in the performance improvements it enables for complete manipulation. If manipulation is performed via learned policies, such policies are likely to use raw sensor data as input (in our case, raw readings from the receiver LEDs), decreasing the importance of any intermediate representations. However, training such a policy exceeds the scope of this paper, which focuses on the design of the sensor itself.
Therefore, in order to quantify the standalone performance of our sensor, we test the ability to infer ground truth wrenches applied to the robotic finger from the raw sensor data. To that end, we mount our sensor on a commercial ATI SI-130-10 to obtain ground truth force and torque measurements. We then trained a simple machine learning model to predict applied force and torque based on our sensor data. We describe the experimental approach and results in this section.
@@ V-A Data Collection
To collect the dataset, we mounted a 3D-printed finger-like shape on top of our sensor, as shown on the right in Figure 1. We made contact with it from all directions along its circumference. Each data file contains three contact events recorded within a 15-second span. To ensure force and torque were decoupled, we applied a range of forces at different heights. The dataset primarily captures forces between 0–5 N and torques between 0–0.2 Nm along each axis, with a maximum force magnitude limited to under 10 N.
We selected 10 contact locations, distributed evenly around the circumference of the finger. Contacts were applied by pressing against the printed finger with our fingertip for 2 and 5 seconds. For the training set, we collected data for 15 different contacts at each location. These contacts were grouped into sets of three, each recorded at different heights—top, middle, and bottom of the finger. Each set focused on a specific force range to ensure comprehensive coverage. The test sets are collected with the same procedure, but included only 1 contact per surface for each contact position. The test set was randomly split across the height groupings to be representative of the different torque-force pairings.

### v3

@@ Abstract
In this paper, we introduce a sensor designed for integration into robotic fingers, where it can provide information on the displacements induced by external contacts. Our sensor uses LEDs to sense the displacement between two plates connected by a transparent elastomer; when a force is applied to the finger, the elastomer displaces and the LED signals change. We show that using LEDs as both light emitters and receivers in this context provides high sensitivity, allowing such an emitter and receiver pairs to detect very small displacements. We characterize the standalone performance of the sensor by testing the ability of a supervised learning model to predict complete force and torque data from its raw signals, and obtain a mean error between 0.05 and 0.07 N across the three directions of force applied to the finger. Our method allows for compact packaging (fitting at the base of a finger) with no amplification electronics, low cost manufacturing, easy integration into a complete hand, and high overload shear forces and bending torques, suggesting future applicability to complete manipulation tasks.
@@
Contact Sensing, Force Sensing, Sensing for Manipulation, Robot Hands
@@ I Introduction
The most direct method for sensing such forces and torques is to add an off-the-shelf six-axis force and torque (F/T) sensor into the base of the finger. However, such F/T sensors are primarily developed with industrial applications in mind [6, 7, 8], leading to limitations when being used for dexterous manipulation. Manipulation often benefits from compliance [9], whereas most F/T sensors are optimized to be as stiff as possible, thus requiring precision engineering and leading to a high cost. Industrial applications also tend to favor a higher resolution than is needed for manipulation, but compromise by limiting the overload protection [6]. This leads to sensors that are easily damaged in the uncertain environments that robotics need to interact with. Finally, off-the-shelf F/T sensors can be difficult to package and integrate into the hand. Attempts to solve this problem by designing sensors to be placed within the fingertips [10, 11, 12] limit the geometry of the finger and can also prevent the inclusion of other modalities within the finger.
@@
In this paper, we present the design of a displacement sensor intended to provide information about the net forces acting on a robot finger in a manner that addresses the needs of robotic manipulation. First, such a sensor must be easy to manufacture using low-cost components and processes, so that it is affordable and practical to include multiple in a multi-fingered hand. Second, it must have the size and profile suited for the base of a fingertip, which will increase ease of integration into a hand. This also frees up the fingertip design for alternate sensors, or allows more flexibility in design of a finger. Third, some amount of compliance could be beneficial to manipulation, and thus there is no need to avoid compliance at all cost. Finally, the ideal sensor would have a Signal to Noise Ratio (SNR) that allows it to be useful for manipulation without sacrificing overload protection.
@@
Figure 1: Our LED-based sensor, standalone (left) and mounted at the base of a robot finger (right). Our sensor uses LEDs mounted on two plates connected by a transparent elastomer to sense displacement between the plates that corresponds to the applied forces and torques. This method allows for a compact, fully integrated package at a low manufacturing cost.
@@
A key feature of our work is that we demonstrate that using LEDs as both emitters and receivers confers a high signal-to-noise ratio, significantly higher than using photodiodes. Combined with the small LED form factor, this allows us to sense very small displacements in a compact package. Using an off-the-shelf F/T sensor as ground truth, we show that our sensor has an observed resolution of .06 N and 2.6 N-mm in a compact, low-cost package that is compliant and easy to integrate in robot fingers as it requires no external amplification. To the best of our knowledge, this is the first time that LEDs have been integrated into a 6-DOF flexure, as both emitters and receivers, thereby allowing for full contact information in a compact and complaint robot finger. Our sensor is designed to be integrated at the base of the finger, which allows for sensing coverage of the entire finger and the ability to integrate additional sensors. We hope that this direction can facilitate truly ubiquitous sensing integrated in robot fingers, with future applications in manipulation.
II Related Work
Many tactile sensors use LEDs due to their compact packaging and wide range of emission cones [15, 16, 17, 18, 19]. The LEDs in these sensors are embedded in a translucent elastomer, so that their emission cone shines on a receiving device. This receiver is often a photodiode [16, 18] or camera [17, 19, 20], but can also be an LED [21]. Regardless of the receiving device, when a force is applied to the surface of these sensors, the elastomer deforms and the emitter and receiver will displace. Using either a data-driven approach or deriving an analytical relationship, the change in electrical signal from this deformation can then be mapped to the contact applied. These tactile sensors are not able to cover the full surface of a robot finger, and therefore do not provide full sensing coverage for all possible contacts. To detect the full contact, others have tried several modalities including strain-based transducers [22], capacitance [23], optical [24], and magnetic [12]. Although most of these sensors are too large to integrate in a robot finger, there are a few novel F/T sensors that are built particularly for manipulation.
@@ III Sensor Design
Figure 2: A diagram of our LED-to-LED sensor concept. We align two PCBs with LEDs directly above each other. We attach these circuit boards to rigid plates, we then suspend both of these plates in a transparent, elastic material. Assuming the bottom plate is fixed, a force applied to the top plate will displace the plate and change how to emitter LED shines on the receiving LEDs. We can then map this change to the contact wrench that induced the displacement.
@@ III-A Concept
The conceptual diagram of our sensor is shown in Fig. 2. It consists of two rigid plates connected by an elastomer layer that allows 6-DOF displacements. The electronics fixed to each rigid plate consist of a circuit boards with emitter LEDs and clusters of receiving LEDs. At the neutral position, the emitter LED is directly above the center of the receiving cluster. When an external force is applied, the plates will displace relative to each other. This displacement will result in a change in signal for the receiver LEDs due to the relative motion of the corresponding emitter LED. We can then map this change in signal back to the contact wrench applied.
@@ III-B Analysis of Single Emitter-Receiver Pair
Using LEDs as both a receiver and emitter gives us several advantages over the other modalities mentioned in Table I, as light can be sensed at very high frequencies with minimal electrical noise [25]. There are several devices we could use as our receiver, but we opted for LEDs based on three factors: 1) LEDs have a smaller form factor than most photodiodes in the price range. 2) LEDs are optimized for operation in their wavelength range, making it easier to design features to minimize external interference. Adding an opaque wrapping to the outside of our sensor can further prevent interference from external light. 3) LEDs have a smaller viewing angle, which is useful for increasing sensitivity given our size constraints.
Figure 3: To analyze single emitter-receiver pairs, we used a precise linear motor to displace an emitting LED while recording the response from both a receiving LED and a photodiode. Left: sharper cone response of the LED compared to the photodiode. Right: zooming in on response across a .1mm displacement, we see significantly higher SNR when using the LED as receiver versus the photodiode.
We first verified the validity of LEDs as both emitters and receivers, and compared to the alternative of using photodiodes on the receiving end. To achieve this, we measured the response to relative horizontal displacements of a single emitter-received pair separated by an air gap, where the receiver consisted of either an LED or a photodiode. The results, plotting receiver signal vs. relative horizontal displacement of the emitter, are shown in Fig. 3. We observe that the LED receiver exhibits much higher sensitivity compared to its photodiode counterpart. In fact, the LED receiver can identify displacements as small as 0.01 mm, whereas the corresponding signal changes for a photodiode receiver are below the noise level. Based on this result, we proceed with LED-LED pairs for the overall sensor design.
Figure 4: Our complete sensor is composed of two custom PCBs that both contain LED emitters and clusters of LED receivers. The PCBs are fastened to rigid, 3-D printed plates. The plates are then fixed above each other in a mold that we then fill with polydimethylsiloxane (PDMS). We use a single JST connector to communicate between the top and bottom board, and an additional JST connector on the bottom board to communicate with external electronics. The rigid plates are fitted with threaded holes to mount external hardware. Our sensor is 27 mm in diameter, 20mm in height, and weighs 16.4g.
@@ III-C Complete Sensor Design
With the core concept of using identical LEDs as emitters and receivers verified, we implemented a complete design for the sensor. As we wanted to mount this at the base of a finger, we constrained our design to a one inch (25.4 mm) diameter. The goal was to maximize the amount of data signals within the space. Based on our LED analysis, we used a compact, infrared LED (SFH 4056) with an emission cone of $20^{\circ}$. The primary size constraint became the operational amplifiers, which meant we could support a maximum of 12 signals per board. As shown in the left of Fig. 4, we divided the signals into clusters of four receivers, so that each board had three clusters, for a total of 24 signals broken across 12 clusters. We then suspended the PCB boards above each other and attached them with an elastic flexure.
We fastened the bottom and top PCB boards, as shown in Fig. 4, each to a rigid, 3-D printed plate. The plates each have three M2 inserts used to mount to external hardware. We used these inserts to attach these rigid plates to a mold that we then filled with an uncured mixture of polydimethylsiloxane (PDMS). We cured the filled mold at $70^{\circ}C$ for 8 hours, which gave us the sensor shown in Fig. 4. The overall dimensions of the sensor are 20.82mm in height and 27mm in diameter, just above our target diameter. We then covered the sensor with a thin, opaque rubber shrink-wrap tube. This cover inhibits interference from external light sources and adds to the sensor robustness. We cured the PDMS at a 10:1 ratio of base to curing agent, as we observed this best achieved the desired sensor stiffness and surface finish. Moreover, this ratio produced the most consistent behavior across multiple sensors.
@@
Figure 5: Mechanical hysteresis testing over three consecutive loading and unloading cycles. We fixed our sensor on a reference F/T sensor and used a linear probe to displace the top plate by 1 mm.
Figure 6: Mechanical overload forces and torques for our sensor. We built five sensors and then used an ADMET Instron machine to apply a horizontal force until the sensor failed. We tested our sensor in five major failure modes - shear, bending, torsion, tension, compression. All tests were done with the opaque, shrink wrapped rubber cover on.
@@ IV Mechanical Characterization
A core design objective of this sensor is robustness against large displacements induced by high contact forces. We tested our sensor to failure in five failure modes - shear, bending, torsion, tension, and compression. We note that, since our sensor is intended to be mounted at the base of a finger, as shown in Fig. 1, contact forces will mostly cause shear and bending of our sensor. Additionally, all tests on our sensor were done with a shrink wrapped neoprene cover, as shown on the finger in Fig. 1.
Results are shown in Fig. 6. In particular, we note that our sensor withstands shear forces of up to 110 N and bending torques of up to 1.62 N*m. The overload forces for torsion and compression are also high. Our sensor is weakest in tension, which is not a major concern as a robot finger would likely not experience large tension forces.
PDMS is a hysteric material, so we expect our 6-DOF flexure to display some time-dependent non-linearity [15, 26]. In Fig. 5, we plot the results of cyclic loading and unloading tests for our sensor. We used a linear probe to horizontally displace the top plate of our sensor while it was fixed to a reference, off-the-shelf F/T sensor. Displacing the top plate by 1 mm three consecutive times, we observe hysteresis that is relatively consistent over the cycles. While hysteresis is a common effect for sensors that use PDMS as an elastic flexure [26], it is also problematic for mapping the electrical signals to the force applied, and often displays behavior that is unique to each sensor. To design a more robust and consistent sensor, we intend to investigate a variety of different flexures, including other clear polymers as well as rigid alternatives such as metal or plastic flexures. Additionally, future work can aim to address hysteresis from a software perspective by adding time-history information to the model processing the raw data.
@@ V Performance
Ultimately, we believe that the true value of a sensor designed for robot fingers lies in the performance improvements it enables for complete manipulation. If manipulation is performed via learned policies, such policies are likely to use raw sensor data as input (in our case, raw readings from the receiver LEDs), decreasing the importance of any intermediate representations. However, training such a policy exceeds the scope of this paper, which focuses on the design of the sensor itself. Therefore, in this section, we aim to quantify the performance of our design as a standalone sensor.
Since we are likely to train policies directly on raw signals produced by contacts on this sensor, we did not formulate an analytical model relating the change in signal to the displacement of the PCBs and the force applied. Instead, in order to quantify the standalone performance of our sensor, we test the ability to infer ground truth wrenches applied to the robotic finger from the raw sensor data. To that end, we mount our sensor on a commercial ATI SI-130-10 to obtain ground truth force and torque measurements. We then trained a simple machine learning model to predict applied force and torque based on our sensor data. We describe the experimental approach and results in this section.
@@ V-A Data Collection
To collect the dataset, we mounted a 3D-printed finger-like shape on top of our sensor and made contact with it from all directions along its circumference (Figure 1) while it was mounted on a reference FT sensor. Each data file contains three contact events recorded within a 15-second span. To ensure force and torque were decoupled, we applied a range of forces at different heights. The dataset primarily captures forces between 0–5 N and torques between 0–0.2 Nm along each axis, with a maximum force magnitude limited to under 10 N.
We selected ten contact locations, distributed evenly around the circumference of the finger. Contacts were applied by pressing against the printed finger with our fingertip for two and five seconds. For the training set, we collected data for 15 different contacts at each location. These contacts were grouped into sets of three, each recorded at different heights—top, middle, and bottom of the finger. Each set focused on a specific force range to ensure comprehensive coverage. The test sets are collected with the same procedure, but included only one contact per surface for each contact position. The test set was randomly split across the height groupings to be representative of the different torque-force pairings.